### v3

@@ III Result
Figure 5: Evolution and adoption of SLAM techniques in BVI navigation systems over time.
@@ III-B1 Advantages
The integration of different types of sensors and technologies with SLAM significantly expands the scope of its application. By integrating techniques such as object detection algorithms or combining RGB-D and IMU sensor data, SLAM systems achieve a multilayered perception of the environment [64, 84]. Sensor fusion enhances a system’s ability to detect and classify objects, accurately navigate, and handle dynamic elements within the environment, thereby offering a more holistic assistive solution [83, 89].
The cost-effectiveness of SLAM-based solutions attributed to their reliance on widely available low-cost sensors makes this technology particularly interesting. Systems employing monocular cameras or wearable RGB-D cameras exemplify how SLAM can be implemented in a cost-effective manner without compromising functionality, thus making advanced navigational aids accessible to more users [69, 85]. The advantages of SLAM, derived from the literature, are listed in Table XI.
TABLE XI: Key features of SLAM that benefit visually impaired navigation.
@@
Accurate localization
SLAM provides precise positioning within an environment, essential for effective navigation and obstacle avoidance.
[56, 57, 58, 62, 64, 67, 68, 69, 70, 73, 74, 75, 76, 77, 78, 80, 82, 83, 85, 88, 89, 90, 91, 92, 93, 95, 97, 98, 99, 102, 105, 106, 107, 108, 109, 110, 112]
Environmental mapping
SLAM constructs maps of surroundings, enabling spatial awareness for navigation.
[57, 58, 59, 60, 61, 64, 67, 69, 70, 72, 73, 74, 75, 76, 77, 79, 83, 84, 87, 88, 89, 90, 92, 93, 96, 97, 100, 104, 107, 108, 109, 111, 112]
Map reuse
Previously created maps can be reused to enhance the efficiency and reduce the need for constant remapping.
[57, 75, 76, 89, 93, 109]
Loop closing
SLAM corrects trajectory drifts by recognizing previously visited locations, thereby improving long-term accuracy.
[57, 63, 84]
Semantic mapping
SLAM integrates contextual information into maps, thereby enhancing the understanding of the environment and its elements.
[66, 84, 100]
Object localization
SLAM identifies and positions objects within the environment, facilitating interaction and navigation around the obstacles.
[72, 85, 88, 91]
Dense navigation maps
SLAM generates detailed maps that provide rich environmental data that are crucial for complex navigation tasks.
[64, 66, 89]
Incremental map updating
SLAM continuously updates maps with new information, ensuring that they remain accurate and current.
[64, 69, 89]
Integration with other technologies
SLAM can be combined with other technologies and algorithms to enhance functionality and performance.
[64, 68, 72, 83, 84, 85, 89, 92, 96, 98]
Integration of sensors
SLAM utilizes a variety of sensors to enrich the environmental perception and mapping accuracy.
[58, 79, 84, 91]
Cost-effectiveness and accessibility
SLAM’s reliance on commonly available sensors makes it an affordable solution for widespread use.
[66, 69, 76, 83, 85]
Ground-truth trajectory
SLAM delivers accurate path tracking, aiding the development of reliable navigation instructions.
[63, 65, 76]
Integration with 2D map
SLAM data can be integrated with 2D maps to enhance the navigation accuracy and functionality.
[63, 70, 75, 77, 80, 98]
Dynamic environment handling
SLAM can be adapted to changes within environment, maintaining reliable navigation in the presence of moving obstacles.
[73, 90, 109]
Re-localization
SLAM can quickly regain accurate positioning after temporary tracking loss, ensuring continuous and reliable navigation.
[57]
@@ III-B2 Limitations
A notable challenge is the computational complexity and the associated demand for system resources. The implementation of advanced SLAM algorithms and the integration of deep learning frameworks for semantic understanding introduce significant computational overhead [57, 72]. This complexity can compromise the real-time performance, which is crucial for assistive navigation. The need for appropriate hardware to process high-resolution data further underscores this limitation, potentially restricting the deployment of SLAM-based systems [88, 89].
The effectiveness of SLAM is dependent on its environmental characteristics. Accurate mapping and localization depend on the presence of distinct geometric features. In environments lacking such features or dynamically changing settings, SLAM systems may struggle to maintain accurate localization, thereby leading to navigation errors [77, 91]. This limitation is particularly evident in feature-poor areas such as long corridors or spaces with uniform surfaces, where loss of localization can occur [62].
Another critical limitation is the dependency on initial data or pre-existing maps. Some SLAM systems require sighted individuals to pre-map the environment, which can limit the flexibility and immediate usability of unmapped or altered spaces [75, 76]. This reliance on prior mapping can be a significant hurdle for deploying SLAM-based navigation aids in diverse and changing environments [82, 96].
@@
Some SLAM applications require external calibration or setup, such as placement of calibration boards in specific environments. This requirement can limit the spontaneity and ease of use of SLAM-based navigation aids because it imposes additional constraints [105]. Table XII outlines the overall limitations of SLAM, derived from the publications under review.
TABLE XII: Limitations of SLAM for visually impaired navigation.
@@
Complexity and computational requirements
SLAM’s advanced algorithms can demand significant computational power, impacting real-time performance and efficiency.
[57, 72, 79, 88, 89, 91]
Dependency on environmental features
The accuracy of SLAM depends on the presence of distinct environmental features, which limits its effectiveness in feature-poor or dynamically changing environments.
[77, 91, 111]
Dependence on initial data/prior maps
Some SLAM systems require pre-mapped environments or initial data setups by sighted individuals, thereby reducing the flexibility in unmapped or altered spaces.
[75, 76, 82, 89, 96]
Loss of localization in feature-poor areas
SLAM may experience frequent localization losses in areas lacking sufficient feature points, such as blank corridors or plain walls, thereby compromising navigation reliability.
[62]
Dependency on external calibration
The need for external calibration in certain SLAM applications can limit their spontaneity and practicality in unprepared environments.
[105]
Drifting error
Accumulating drifting errors in SLAM can reduce the long-term accuracy, leading to potential navigation inaccuracies and user disorientation.
[56]
Lack of dense navigation maps
The inability of certain SLAM systems to generate detailed dense maps can restrict their effectiveness in providing a comprehensive navigation guidance.
[64]
Vulnerability in dynamic environments
SLAM systems may struggle to adapt to dynamic environments with moving obstacles, thereby posing navigation challenges and safety risks.
[68]
@@
Optimal pathfinding, perception of the surroundings, and obstacle avoidance are crucial for navigation. The precise localization of a visually impaired user within the environment is essential for these functions to operate effectively. Our surveyed papers addressed the localization and mapping problems using various techniques. Some of these studies explored other navigation-related challenges. We categorized these challenges into two groups: those relevant to environmental complexities, and those related to the sensors used to receive environmental data. Dynamic obstacles and crowded spaces constitute the challenges in the first group, whereas challenges related to changing lighting conditions and the rapid motion of users that results in motion blur fall into the second group. Table XIII lists studies that investigated these challenges through the integration of SLAM with other approaches. In the following section, we discuss the studies that address these challenges.
TABLE XIII: Challenges addressed in reviewed studies using SLAM techniques

### v4

@@ III-A1 Approaches
Figure 4: Evolution and adoption of localization and mapping techniques in BVI navigation systems over time.
III-A2 Sensor type
The sensors employed in SLAM solutions for BVI navigation are diverse and include various types of cameras, LiDAR, IMU, and other specialized sensors. As shown in Figure 5, we categorized the sensors into three types: cameras, LiDAR, and other sensors.
TABLE XVII: Comparison of core technical features for localization and mapping techniques, with a specific focus on sensor types, computing resources, and whether machine learning-based methods are employed for localization and mapping tasks - Part I.
Ref. Sensor type Computing Resource Localization & Mapping Technique ML-Based Localization and Mapping
[56] Camera Smartphone Native AI library for iOS devices Not explicitly mentioned
[57] RGB-D camera Remote server ORB-SLAM2 $\bullet$
[58] A depth, an RGB, and four gray scale cameras, an IMU Hololens2 device, GPU Iterative Closest Point (ICP) for camera pose estimation, Kinect Fusion algorithm for real-time 3D reconstruction Not explicitly mentioned
[59] Wheel encoder and LiDAR Notebook PC Cartographer $\bullet$
[60] RGB camera Cloud server and Nvidia Jetson AGX Xavier Visual place recognition, weighted averaging, and perspective-n-point (PnP) for localization, OpenVSLAM and Colmap to generate a topometric map NetVLAD for global descriptors and SuperPoint for local descriptors
[61] LiDAR Not mentioned Cartographer to build SLAM maps on each floor $\bullet$
[62] RGB-D camera Jetson AGX Xavier ORB-SLAM2 for localization; down-sampling, octomap, and 2D occupancy grid mapping for an accurate dense mapping $\bullet$
[63] IMU, stereo and IR (depth) cameras Cloud server ORB-SLAM3 $\bullet$
[64] RGB-D camera Raspberry Pi ORB-SLAM $\bullet$
[65] Monocular RGB Not mentioned ORB-SLAM3 to obtain ground-truth camera trajectory $\bullet$
[66] RGB-D camera High performance portable processor, cloud server Semantic visual SLAM based on ORB feature to generate sparse, dense, and semantic maps ENet for pixel-level semantic segmentation
[67] LiDAR Raspberry Pi BreezySLAM $\bullet$
[68] LiDAR Intel NUC computer GMapping for mapping and estimating the destination’s location $\bullet$
[69] RGB-D camera Raspberry Pi3 B+ Hector SLAM for building the environmental map and locating the user on the map $\bullet$
[70] RGB-D Camera, IMU UP Board computer A VIO system developed based on VINS-Mono $\bullet$
[71] RGB-D-IR camera A portable computer, Nvidia Jetson TX2 Hierarchical visual localization pipeline with deep descriptors, geometric verification, and sequence matching. NetVLAD and Dense Desc for advanced descriptor extraction.
[72] RGB-D camera Odroid XU3 board, remote server Off-the-shelf algorithm [120] for 3D indoor mapping; a two-stream ConvNet for topological localization. ConvNet replaces BoW for semantic info; Inception-v3 enhances object recognition, aiding localization.
[73] Two monocular cameras Not mentioned Followed a similar structure as ORB-SLAM2. $\bullet$
[74] RGB-D camera, IMU A smartphone with Qualcomm Snapdragon 820 CPU 2.0 GHz Vins-mono for indoor localization; ORB-SLAM2 and Vins-mono for building a indoor map repository. $\bullet$
[75] Smartphone’s camera A HUAWEI P20 smartphone with Kirin 970 CPU Roberto Lopez Mendez ARCore SLAM as the base for visual odometry and area learning Not explicitly mentioned
[76] Fisheye and depth camera An embedded CPU board ORB-SLAM2 for the building of the virtual-blind-road (offline by a sighted person) and the localization (online) $\bullet$
[77] SwissRanger SR4000 3D camera A Lenovo ThinkPad T430 laptop 2-step graph SLAM $\bullet$
[78] SwissRanger SR4000 camera, IMU Up Board computer PAVIO: Fusing visual, inertial, and plane features for robust SLAM localization and mapping. $\bullet$
[79] Encoder, IMU, laser distance sensor Raspberry Pi 3 Model B and B+ Gmapping for building a 2D occupancy grid map on the environment $\bullet$
[80] iPhone 11 Pro camera and IMU iPhone 11 Pro ARKit VIO for relative movements; ARKit mapping for semantic labels; Monte Carlo localization YOLOv2 for object detection to facilitate effective localization
TABLE XVIII: Comparison of core technical features for localization and mapping techniques, with a specific focus on sensor types, computing resources, and whether machine learning-based methods are employed for localization and mapping tasks - Part II.
Ref. Sensor type Computing Resource Localization & Mapping Technique ML-Based Localization and Mapping
[82] RGB-D camera, compass sensor Jetson Xavier AGX, NVidia FastSLAM for mapping, ORB-SLAM2 for local pose estimation in the pre-built map An extra thread aligns predicted semantics with key features.
[83] LiDAR, ultrasonic sensor, Raspberry pi camera Raspberry Pi3 B+ Hector SLAM for constructing a 2D-map of the environments and localization $\bullet$
[84] RGB-D camera, IMU Nvidia Jetson AGX Xavier microprocessor RTAB-Map for semantic point cloud generation and global localization MobileNetV2 with PPM for constructing semantic point cloud
[85] Monocular camera Nvidia Jetson Xavier NX Developer kit Visual SLAM for user movement estimation and stationary object localization $\bullet$
[87] LiDAR, IMU Nvidia RTX 3080 graphic board Cartographer for constructing a LiDAR map $\bullet$
[88] RGB-D camera, IMU Dell G15 laptop with an RTX 3060 GPU SLAM implementation by RealSense for creating an initial 2D occupancy grid and estimating user pose $\bullet$
[89] RBG-D camera Uzel US-M5422 edge server, Raspberry Pi 4B Improved ORB-SLAM for generating a dense navigation map and real-time positioning $\bullet$
[90] RGB-D camera Laptop Dynamic-SLAM based on ORB-SLAM2 for estimating user’s ego-pose and building a static feature point map Non-Prior dynamic object detection preceding local mapping
[91] RGB-D camera, IMU Google Pixel 3 smartphone RGBD-VIO for mapping and accurately estimating the device’s pose $\bullet$
[92] RealSense D435i camera Remote server based on Intel i7-8700 CPU, nvidia GTX1080 GPU ORB-SLAM2 for environmental mapping and positioning of the user $\bullet$
[93] Google glasses camera GPU server OpenVSLAM for mapping and locating user’s indoor positions $\bullet$
[95] LiDAR, IMU Laptop Cartographer for estimating the current location and direction of a user $\bullet$
[96] RGB-D camera, IMU Not mentioned ORB-SLAM2 for mapping and localizing the user $\bullet$
[97] RGB-D camera Nvidia Jetson AGX Xavier processor OpenVSLAM for robust mapping and localization in real time. $\bullet$
[98] iPhone 8’s IMU and rear-facing camera iPhone 8 Visual-Inertial Odometry with sign recognition and geometric constraints. Not explicitly mentioned
[99] RGB-D camera, IMU UP Board computer Visual-Inertial Odometry for 3D mapping and pose estimation $\bullet$
[100] RGB-D, LiDAR Laptop Semantic SLAM based on a 2D SLAM technique for determining the corridor area and mapping to a semantic map YOLOv3 for landmark detection, Places365 for place recognition
[101] Vuze+ camera Not mentioned ORB-SLAM2 for generating 3D positions and 2D connectivity graph from Vuze+ footage $\bullet$
[102] Monocular camera Intel i7 processor Non-filter based visual SLAM using trained objects as landmarks for localization ACF detector for object detection to identify trained objects of interest for localization.
[103] ZED camera Nvidia Jetson TX2 ZED camera’s SLAM for navigating toward an aisle in a grocery store. Not mentioned
[104] Stereo camera, laser Nvidia Jetson TX2, Raspberry Pi, and Arduino. RTAB-MAP for obtaining fine-grained mapping of the 3D spatial world. $\bullet$
[105] Monocular camera Not mentioned Visual odometry for user localization during street crossings $\bullet$
[106] 3D time-of-flight camera , IMU UP Board computer Plane-aided visual-inertial odometry (PAVIO) for pose estimation of an robotic navigation aid $\bullet$
[107] Wide-angle lens camera, gyroscope, accelerometers, and infrared sensor on Google Tango. Google Tango Google Tango’s built-in SLAM for mapping the environment and localizing the user. $\bullet$
[108] Stereo camera Cloud server Visual SLAM for mapping the environment and localizing the user. $\bullet$
TABLE XIX: Comparison of core technical features for localization and mapping techniques, with a specific focus on sensor types, computing resources, and whether machine learning-based methods are employed for localization and mapping tasks - Part III.
Ref. Sensor type Computing Resource Localization & Mapping Technique ML-Based Localization and Mapping
[109] Monocular camera A single CPU LSD-SLAM for estimating the user’s position and constructing a 3D environmental map $\bullet$
[110] Camera, IMU, LiDAR Smartphone visual-inertial SLAM based pose estimation and 2D scene-graph map construction using iOS ARKit. Not explicitly mentioned
[111] Kinect camera Not mentioned LSD-SLAM for assessing the user’s location and building an environmental map $\bullet$
[112] Laser Range Finder (LRF) sensor PC Hector SLAM for map building and odometry calculation $\bullet$
Figure 5: Overview of sensor types in studies under review
Camera The common use of visual sensors in SLAM techniques can be attributed to advances in computer vision and image processing, which enhance navigation capabilities by providing rich environmental information. This makes visual-sensor-based SLAM techniques the most commonly used in the implementation of assistive technologies for the BVI people, offering a cost-effective, versatile, and accurate solution for real-time navigation and spatial awareness. The literature under review used the following types of cameras: RGB, RGB-D, stereo, monocular, and other specialized cameras.
RGB cameras are widely used due to their ability to capture rich color information, which is beneficial for visual odometry and object recognition. They are cost effective and widely available, making them a popular choice for the development of accessible navigation aids.
RGB-D cameras provide both color and depth information, enabling more accurate mapping and localization. Depth information helps in understanding the 3D structure of the environment.
Stereo cameras also provide depth perception through two slightly offset lenses that simulate binocular vision. They are effective in capturing detailed depth information and are useful in applications where precise depth estimation is required.
Monocular cameras are simpler than stereo and RGB-D cameras. They rely on visual odometry and other techniques to estimate depth and motion, making them lightweight and suitable for mobile applications.
Specialized cameras, including fisheye, 3D time-of-flight, and wide-angle cameras, provide specialized capabilities, such as a wide field of view or precise depth measurement, which can enhance the SLAM performance in specific scenarios.
LiDAR LiDAR sensors are highly accurate in measuring distances and are effective in creating detailed 3D maps of the environment. Studies use LiDAR alone to build a map of the environment or in combination with other sensors such as IMU and cameras to enhance the robustness and accuracy of SLAM systems.
Other sensors Various studies combined different types of sensors to leverage the strengths of each type and provide more robust and reliable navigation solutions. For example, integrating an IMU with a camera helps achieve better motion tracking and stability. This trend towards integrating multiple sensors highlights increasing efforts to enhance the robustness and reliability of SLAM solutions.
The reviewed papers show a clear preference for RGB-D cameras, indicating their effectiveness in providing both the visual and depth information necessary for accurate SLAM applications. The use of LiDAR is important in applications that require precise mapping. Over the years, there has been a noticeable trend towards integrating multiple sensors and combining their strengths to achieve more robust and reliable SLAM solutions for visually impaired navigation. The integration of machine learning-based techniques with SLAM systems is particularly prevalent in solutions that utilize RGB-D cameras. This highlights the effectiveness of combining this type of data with advanced AI algorithms. This trend is likely to continue as technology advances, offering more sophisticated and adaptable solutions in complex and dynamic environments.
III-A3 Computing resource
To process data and run localization and mapping algorithms, the reviewed studies adopted two classes of computational resources: local and remote. Local computations are performed in situ on devices, such as smartphones, tablets, laptops, portable microcontrollers, and UP board computers. In some cases, algorithms were applied on PCs. Table XX categorizes the computing resources used in the reviewed studies for localization and mapping tasks. Information regarding the computing resources for the entire navigation assistive system is detailed in Tables XXXV-XXXVII
TABLE XX: Categorization of computing resources used for localization and mapping tasks.
Computing resource Description References
Remote/Cloud servers Used for high-computation tasks, leveraging the power of remote resources. [72, 60, 57, 63, 66, 92, 93, 108]
Smartphones/Tablets Common in applications prioritizing portability and accessibility. [107, 56, 75, 80, 91, 98, 110, 74]
Nvidia Jetson For balancing computational power and portability. [82, 104, 60, 62, 97, 71, 84, 85, 87, 103]
Raspberry Pi Chosen for its affordability and sufficient power. [104, 64, 67, 69, 79, 83, 89]
Laptops and PCs Employed for tasks needing robust computational capabilities and flexibility in hardware. [59, 90, 88, 100, 77, 95, 112]
UP Board computer Used for handling intensive computation tasks while maintaining a compact form factor. [70, 78, 99, 106]
Other specific systems Includes a variety of specific embedded solutions tailored to the requirements of each study. [76, 102, 104, 109, 68, 72, 58, 89]
Local Computing Resources
Smartphones: Smartphone are widely used as communication gadgets, and their technology continues to grow to the point that it is possible for smartphones to implement functional navigation systems. Because smartphones integrate diverse sensors such as IMU, GPS, and cameras, they can be used as a convenient tool for collecting environmental information. In addition, their computational power can be exploited to perform various navigation operations. For example, the system proposed by [74] implemented all algorithms relevant to data acquisition, ground segmentation, moving direction search, global path planning, indoor and outdoor localization, and object detection on a smartphone and achieved real-time performance. Without an additional depth sensor, [75] took advantage of an ARCore-supported smartphone to track pose and to build a map of the surroundings in real time. However, despite the significant advantages of smartphones, such as their small size, low weight, easy portability, and low cost, their computing power is not sufficient for some approaches.
Laptops and PCs: Some of the reviewed approaches perform all or part of the required calculations locally on a portable computer, such as a laptop. Despite higher computing power compared to a smartphone, and greater security compared to remote computational resources, the laptop’s heavier weight and large size are considered major disadvantages, especially during long trips. PCs provide even higher computing power, which is essential for complex SLAM operations; however, they lack portability.
Embedded systems and microcontrollers: Embedded systems such as Nvidia Jetson boards, Raspberry Pi, and UP boards provide a balance between computational power and portability. They are commonly used in the reviewed studies for performing SLAM operations locally. For instance, [58] utilized a Hololens2 device with a GPU for real-time 3D reconstruction, whereas [62] used a Jetson AGX Xavier for ORB-SLAM2 and dense mapping. Raspberry Pi devices are also popular due to their low cost and sufficient computing power for many SLAM tasks [69, 67].
Remote Computing Resources
An alternative solution is to transfer all or part of the calculations to the remote computing resources. To reduce local computing costs, [57] adopted an embedded computer and a remote server. In the proposed vision-based assistance system, before transferring the input images to the server, the images were time-stamped and encrypted on an embedded computer. The remote server was equipped with a CPU and GPU to run parallel ORB-SLAM2 and artificial intelligence algorithms for indoor navigation, object detection, face recognition, and scene text recognition. Experiments confirmed that the use of remote servers under a smooth network connection, such as 4G or WiFi, can meet the computational requirements of the proposed system. However, although the high computing power of remote servers is considered a significant advantage, constant Internet access over a secure connection is required. Moreover, the performance of the entire system would be affected by the network condition.
@@ III-B1 Advantages
The integration of different types of sensors and technologies with SLAM significantly expands the scope of its application. By integrating techniques such as object detection algorithms or combining RGB-D and IMU sensor data, SLAM systems achieve a multilayered perception of the environment [64, 84]. Sensor fusion enhances a system’s ability to detect and classify objects, accurately navigate, and handle dynamic elements within the environment, thereby offering a more holistic assistive solution [83].
The cost-effectiveness of SLAM-based solutions attributed to their reliance on widely available low-cost sensors makes this technology particularly interesting. Systems employing monocular cameras or wearable RGB-D cameras exemplify how SLAM can be implemented in a cost-effective manner without compromising functionality, thus making advanced navigational aids accessible to more users [69, 85]. The advantages of SLAM, derived from the literature, are listed in Table XXI.
TABLE XXI: Key features of SLAM that benefit visually impaired navigation.
@@
Accurate localization SLAM provides precise positioning within an environment, essential for effective navigation and obstacle avoidance. [56, 57, 58, 62, 64, 67, 68, 69, 70, 73, 74, 75, 76, 77, 78, 82, 83, 85, 88, 89, 90, 91, 92, 93, 95, 97, 98, 99, 102, 106, 107, 108, 109]
Environmental mapping SLAM constructs maps of surroundings, enabling spatial awareness for navigation. [110, 99, 57, 58, 59, 60, 61, 64, 67, 69, 70, 72, 73, 74, 75, 76, 77, 79, 83, 84, 87, 88, 89, 90, 92, 93, 96, 97, 100, 104, 107, 108, 109, 111, 112]
Map reuse Previously created maps can be reused to enhance the efficiency and reduce the need for constant remapping. [57, 75, 76, 93, 109]
Loop closing SLAM corrects trajectory drifts by recognizing previously visited locations, thereby improving long-term accuracy. [57, 63, 84]
Semantic mapping SLAM integrates contextual information into maps, thereby enhancing the understanding of the environment and its elements. [66, 84, 100]
Object localization SLAM identifies and positions objects within the environment, facilitating interaction and navigation around the obstacles. [72, 85, 88, 91]
Dense navigation maps SLAM generates detailed maps that provide rich environmental data that are crucial for complex navigation tasks. [64, 66, 89]
Incremental map updating SLAM continuously updates maps with new information, ensuring that they remain accurate and current. [111, 109, 64, 69, 89]
Integration with other technologies SLAM can be combined with other technologies and algorithms to enhance functionality and performance. [102, 64, 68, 72, 83, 84, 85, 89, 92, 98]
Integration of sensors SLAM utilizes a variety of sensors to enrich the environmental perception and mapping accuracy. [104, 100, 99, 58, 84, 91]
Cost-effectiveness and accessibility SLAM’s reliance on commonly available sensors makes it an affordable solution for widespread use. [66, 69, 76, 83, 85, 57, 60, 67, 73, 74, 75, 80, 98, 102]
Ground-truth trajectory SLAM delivers accurate path tracking, aiding the development of reliable navigation instructions. [63, 65]
Integration with 2D map SLAM data can be integrated with 2D maps to enhance the navigation accuracy and functionality. [100, 99, 63, 70, 75, 77, 80, 98]
Dynamic environment handling SLAM can be adapted to changes within environment, maintaining reliable navigation in the presence of moving obstacles. [111, 104, 100, 73, 90, 109]
Re-localization SLAM can quickly regain accurate positioning after temporary tracking loss, ensuring continuous and reliable navigation. [57]
@@ III-B2 Limitations
A notable challenge is the computational complexity and the associated demand for system resources. The implementation of advanced SLAM algorithms and the integration of deep learning frameworks for semantic understanding introduce significant computational overhead [57, 72]. This complexity can compromise the real-time performance, which is crucial for assistive navigation. The need for appropriate hardware to process high-resolution data further underscores this limitation, potentially restricting the deployment of SLAM-based systems [89].
The effectiveness of SLAM is dependent on its environmental characteristics. Accurate mapping and localization depend on the presence of distinct geometric features. In environments lacking such features or dynamically changing settings, SLAM systems may struggle to maintain accurate localization, thereby leading to navigation errors [77]. This limitation is particularly evident in feature-poor areas such as long corridors or spaces with uniform surfaces, where loss of localization can occur [62].
Another critical limitation is the dependence on initial data or pre-existing maps. Some SLAM systems require sighted individuals to pre-map the environment, which can limit the flexibility and immediate usability of unmapped or altered spaces [75, 76]. This reliance on prior mapping can be a significant hurdle for deploying SLAM-based navigation aids in diverse and changing environments [82, 96].
@@
Some SLAM applications require external calibration or setup, such as placement of calibration boards in specific environments. This requirement can limit the spontaneity and ease of use of SLAM-based navigation aids because it imposes additional constraints [105]. Table XXII outlines the overall limitations of SLAM, derived from the publications under review.
TABLE XXII: Limitations of SLAM for visually impaired navigation.
@@
Complexity and computational requirements SLAM’s advanced algorithms can demand significant computational power, impacting real-time performance and efficiency. [57, 72, 89, 90]
Dependency on environmental features The accuracy of SLAM depends on the presence of distinct environmental features, which limits its effectiveness in feature-poor or dynamically changing environments. [109, 77, 111]
Dependence on initial data/prior maps Some SLAM systems require pre-mapped environments or initial data setups by sighted individuals, thereby reducing the flexibility in unmapped or altered spaces. [112, 109, 75, 76, 82, 96, 84, 93]
Loss of localization in feature-poor areas SLAM may experience frequent localization losses in areas lacking sufficient feature points, such as blank corridors or plain walls, thereby compromising navigation reliability. [62]
Dependency on external calibration The need for external calibration in certain SLAM applications can limit their spontaneity and practicality in unprepared environments. [105]
Drifting error Accumulating drifting errors in SLAM can reduce the long-term accuracy, leading to potential navigation inaccuracies and user disorientation. [56]
Dense maps fail to align with real-world conditions The inability of certain SLAM systems to generate detailed dense maps can restrict their effectiveness in providing a comprehensive navigation guidance. [64]
Vulnerability in dynamic environments SLAM systems may struggle to adapt to dynamic environments with moving obstacles, thereby posing navigation challenges and safety risks. [68]
@@
This section explores various challenging situations addressed by SLAM-based navigation-assistive systems for BVI individuals. We categorized these challenges into two main groups: those relevant to environmental complexities and those related to the sensors used for receiving environmental data. Additionally, we discuss practical challenges and considerations that impact the usability and adoption of these systems.
III-C1 Technical and methodological challenges
Optimal pathfinding, perception of surroundings, and obstacle avoidance are crucial for navigation. Precise localization of a visually impaired user within the environment is essential for the effective operation of these functions. Our surveyed papers addressed the localization and mapping problems using various techniques. Some of these studies explored other navigation-related challenges. We categorized these challenges into two groups: those relevant to environmental complexities, and those related to the sensors used to receive environmental data. Dynamic obstacles and crowded spaces constitute the challenges in the first group, whereas challenges related to changing lighting conditions and the rapid motion of users that results in motion blur fall into the second group. Table XXIII lists studies that investigated these challenges through the integration of SLAM with other approaches. In the following section, we discuss the studies that address these challenges.
TABLE XXIII: Challenges addressed in reviewed studies using SLAM techniques